OpenStudy (anonymous):

What values for 0 (0<0pi) satisfy the equation?

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

No idea...

OpenStudy (anonymous):

\[\theta (0 \le \theta \pi )\]

OpenStudy (agent0smith):

still no idea, makes no sense.

OpenStudy (anonymous):

lol okay then cx

OpenStudy (agent0smith):

Post a screenshot...

OpenStudy (anonymous):

no its okay haha im just trying to get this done :P

OpenStudy (agent0smith):

haha okay